Help will thumbs up.
Exercise 22-13 Glacial Company estimates that variable costs will be 67.1% of sales, and fixed costs wil total $696,000 The selling price of the product is $3.80 Compute the break-even point in (1) units and (2) collars (Round intermediate calculation to 2 decimal places, e-g 52.75 and final answers to O decimal p e.g. 5,275.) (1) Break-even sales (2) Break-even sales units Assuming actual sales are $2,454,000, compute the margin of safety in (1) dollars and (2) as a ratio. (Round ratio to 0 decimal (1) Margin of safety (2) Margin of safety ratio places, e.g. 20.) LINK TO TEXT LINK TO TEXT Question Attempts: 0 of 3 used SAVE FOR LATER SUBMET ANSWER
Answer #1

Contribution margin per unit = (100-67.1)*3.80 = 1.25

Break even point units = 696000/1.25 = 556800 units

Break even point sales = 556800*3.8 = 2115840

2) Margin of safety = 2454000-2115840 = 338160

Margin of safety ratio = 338160*100/2454000 = 14%